# Prediction Instability in Machine Learning Ensembles

### v1

@@ 5 Conclusion
In this note we have analyzed a set of four desirable properties on ensemble information use and prediction stability and shown that at least one must be violated in any machine learning ensemble, establishing a fundamental tradeoff between information use and prediction instability. Any ensemble that uses the insights of multiple models will make choices in a counterintuitive or unstable fashion at least some of the time. Since three of these four properties are fairly easy to satisfy the last–model choice reversal–must fail in many common use cases thus introducing ‘discontinuities’ into the predictions and decisions made by ensembles. Small changes in underlying model scores can lead to large, substantive changes in ensemble predictions and choices; the ensemble can change its prediction or choice even if none of the underlying models do and there is no guarantee that similar feature space inputs will be treated similarly.
@@
We observe that the ECD theorem presented here shows that all ensembles will behave badly some of the time with respect to the prediction stability properties we have analyzed. It seems intuitive that the risk of prediction or choice instability depends on the number of models in the underlying ensemble and on the relationship between the size of the data sample used to train the ensemble and the population it is drawn from. Accordingly we aim to extend the analysis here to investigate asymptotic ensemble behavior when the number of models in the ensemble grows large. We conjecture that the Borda count method of aggregation may satisfy definitions 2.1, 2.2, and 2.3 with a slight weakening of model choice reversal in the asymptotic case of many models. We also conjecture that as the data available for training grows arbitrarily large this will act as an implicit restriction on the diversity of the models in the ensemble and so minimize the risk of prediction instability. Finally, we conjecture that the argument developed here would apply to using a data sampling process like bootstrapping or $k$-fold cross validation to choose among multiple model architectures during training.
Acknowledgements
@@
This work was supported by the exceptional generosity of the PPC Foundation, Inc. Additionally, we would like to thank Ernesto Guerra–Vallejos, RJ Nowling, David Andrzejewski, and the audience of the Medical College of Wisconsin SEAWINDS Data Science Symposium for helpful comments, suggestions, and feedback. Finally, we thank Christopher Taylor and Sheila Ross for their support.
@@
Lemma 5.1.
If $C(\cdot,s)$ satisfies transitivity then:
@@
    $y\in C(\{y,y^{\prime}\},s)$ and $y^{\prime}=C(\{y^{\prime},y^{\prime\prime}\},s)$ implies that $y=C(\{y,y^{\prime\prime}\},s)$;
@@
    $y=C(\{y,y^{\prime}\},s)$ and $y^{\prime}\in C(\{y^{\prime},y^{\prime\prime}\},s)$ implies that $y=C(\{y,y^{\prime\prime}\},s)$;
@@
    $y=C(\{y,y^{\prime}\},s)$ and $y^{\prime}=C(\{y^{\prime},y^{\prime\prime}\},s)$ implies that $y=C(\{y,y^{\prime\prime}\},s)$.
@@
Suppose that $C(\cdot)$ satisfies transitivity. To see part 1) assume $y\in C(\{y,y^{\prime}\},s)$ and $y^{\prime}=C(\{y^{\prime},y^{\prime\prime}\},s)$. Then by transitivity $y\in C(\{y,y^{\prime\prime}\},s)$. Suppose that $y^{\prime\prime}\in C(\{y,y^{\prime\prime}\},s)$. But then by transitivity $y^{\prime\prime}\in C(\{y^{\prime},y^{\prime\prime}\},s)$, a contradiction. To see part 2) assume $y=C(\{y,y^{\prime}\},s)$ and $y^{\prime}\in C(\{y^{\prime},y^{\prime\prime}\},s)$. Then by transitivity $y\in C(\{y,y^{\prime\prime}\},s)$. Suppose that $y^{\prime\prime}\in C(\{y,y^{\prime\prime}\},s)$. But then by transitivity $y^{\prime}\in C(\{y,y^{\prime}\},s)$, a contradiction. To see, part 3) assume that $y=C(\{y,y^{\prime}\},s)$ and $y^{\prime}=C(\{y^{\prime},y^{\prime\prime}\},s)$ but that $y^{\prime\prime}\in C(\{y,y^{\prime\prime}\},s)$. Then by part 1) we have $y^{\prime\prime}\in C(\{y^{\prime},y^{\prime\prime}\},s)$, a contradiction. ∎
Proof of Lemma 3.2.
@@
Sufficiency. Suppose that $C(\cdot)$ is transitive. Take any $s\in S(Y)$, any $Y_{1}^{*},Y_{2}^{*}\subseteq Y$ such that $Y_{1}^{*}\subseteq Y_{2}^{*}$. Suppose that $Y_{1}^{*}\bigcap C(Y_{2}^{*},s)\neq\emptyset$. Take any $y\in C(Y_{1}^{*},s)$; we wish to show that $y\in Y_{1}^{*}\bigcap C(Y_{2}^{*},s)$. Clearly $y\in Y_{1}^{*}$. Suppose that $y\notin C(Y_{2}^{*},s)$. Take any $y^{\prime}\in Y_{1}^{*}\bigcap C(Y_{2}^{*},s)$. Then $y^{\prime}\in C(Y_{2}^{*},s)$ implies that $y^{\prime}\in C(,s)$, and $y\in C(Y_{1}^{*},s)$ implies that $\{y,y^{\prime}\}=C(\{y,y^{\prime}\},s)$. Since $y\notin C(Y_{2}^{*},s)$ it follows that there exists $y^{\prime\prime}\in Y_{2}^{*}$ such that $y^{\prime\prime}=C(\{y,y^{\prime\prime}\},s)$. But then by transitivity and Lemma 5.1 it must be that $y^{\prime\prime}=C(\{y^{\prime},y^{\prime\prime}\},s)$, a contradiction. Next take any $y\in Y_{1}^{*}\bigcap C(Y_{2}^{*},s)$. Then there does not exist $y^{\prime}\in Y_{2}^{*}$ such that $y^{\prime}=C(\{y,y^{\prime}\},s)$. Thus there does not exist $y^{\prime}\in Y_{1}^{*}$ such that $y^{\prime}=C(\{y,y^{\prime}\},s)$ and so $y\in C(Y_{1}^{*},s)$. It follows that $C(Y_{1}^{*},s)=Y_{1}^{*}\bigcap C(Y_{2}^{*},s)$ and so definition 2.1 holds.
@@
Necessity. Suppose that $C(\cdot)$ satisfies definition 2.1 but that transitivity does not hold. Then there exists $y,y^{\prime},y^{\prime\prime}$ such that $y\in C(\{y,y^{\prime}\},s)$ and $y^{\prime}\in C(\{y^{\prime},y^{\prime\prime}\},s)$, yet $y^{\prime\prime}=C(\{y,y^{\prime\prime}\},s)$. There are four cases to consider.
@@
Case 1: $\{y,y^{\prime}\}=C(\{y,y^{\prime}\},s)$, $\{y^{\prime},y^{\prime\prime}\}=C(\{y^{\prime},y^{\prime\prime}\},s)$, and $y^{\prime\prime}=C(\{y,y^{\prime\prime}\},s)$. By definition of $C(\cdot)$ we have $C(\{y,y^{\prime},y^{\prime\prime}\},s)=\{y^{\prime},y^{\prime\prime}\}$. But then we have $\{y,y^{\prime}\}\subseteq\{y,y^{\prime},y^{\prime\prime}\}$ and $\{y,y^{\prime}\}\bigcap C(\{y,y^{\prime},y^{\prime\prime}\},s)\neq\emptyset$ and yet $C(\{y,y^{\prime}\},s)=\{y,y^{\prime}\}\neq y^{\prime}=\{y,y^{\prime}\}\bigcap C(\{y,y^{\prime},y^{\prime\prime}\},s)$, a violation of definition 2.1 and so a contradiction.
@@
Case 2: $\{y,y^{\prime}\}=C(\{y,y^{\prime}\},s)$, $y^{\prime}=C(\{y^{\prime},y^{\prime\prime}\},s)$, and $y^{\prime\prime}=C(\{y,y^{\prime\prime}\},s)$. By definition of $C(\cdot)$ we have $C(\{y,y^{\prime},y^{\prime\prime}\},s)=y^{\prime}$. But then we have $\{y,y^{\prime}\}\subseteq\{y,y^{\prime},y^{\prime\prime}\}$ and $\{y,y^{\prime}\}\bigcap C(\{y,y^{\prime},y^{\prime\prime}\},s)\neq\emptyset$ and yet $C(\{y,y^{\prime}\},s)=\{y,y^{\prime}\}\neq y^{\prime}=\{y,y^{\prime}\}\bigcap C(\{y,y^{\prime},y^{\prime\prime}\},s)$, a violation of definition 2.1 and so a contradiction.
@@
Case 3: $y=C(\{y,y^{\prime}\},s)$, $\{y^{\prime},y^{\prime\prime}\}=C(\{y^{\prime},y^{\prime\prime}\},s)$, and $y^{\prime\prime}=C(\{y,y^{\prime\prime}\},s)$. By definition of $C(\cdot)$ we have $C(\{y,y^{\prime},y^{\prime\prime}\},s)=y^{\prime\prime}$. But then we have $\{y^{\prime},y^{\prime\prime}\}\subseteq\{y,y^{\prime},y^{\prime\prime}\}$ and $\{y^{\prime},y^{\prime\prime}\}\bigcap C(\{y,y^{\prime},y^{\prime\prime}\},s)\neq\emptyset$ and yet $C(\{y^{\prime},y^{\prime\prime}\},s)=\{y^{\prime},y^{\prime\prime}\}\neq y^{\prime\prime}=\{y^{\prime},y^{\prime\prime}\}\bigcap C(\{y,y^{\prime},y^{\prime\prime}\},s)$, a violation of definition 2.1 and so a contradiction.
Case 4: $y=C(\{y,y^{\prime}\},s)$, $y^{\prime}=C(\{y^{\prime},y^{\prime\prime}\},s)$, and $y^{\prime\prime}=C(\{y,y^{\prime\prime}\},s)$. But then $C(\{y,y^{\prime},y^{\prime\prime}\},s)=\emptyset$, a contradiction. ∎
Proof of Lemma 3.3.
@@
It suffices to consider two cases. First suppose that $C(\{y,y^{\prime}\},s)=\{y,y^{\prime}\}$ but that without loss of generality $C(\{y,y^{\prime}\},s^{\prime})=\{y\}$. Then there is an immediate contradiction of model choice reversal, since there does not exist $j$ such that $s_{j}(\cdot,y^{\prime})>s_{j}(\cdot,y)$ and $s_{j}^{\prime}(\cdot,y^{\prime})<s_{j}^{\prime}(\cdot,y)$ and so the claim holds. Next suppose without loss of generality that $C(\{y,y^{\prime}\},s)=\{y\}$ but that $y^{\prime}\in C(\{y,y^{\prime}\},s^{\prime})$. But then again there is an immediate contradiction of model choice reversal, since there does not exist $j$ such that $s_{j}(\cdot,y^{\prime})<s_{j}(\cdot,y)$ and $s_{j}^{\prime}(\cdot,y^{\prime})>s_{j}^{\prime}(\cdot,y)$ and so the claim holds. Thus $C(\{y,y^{\prime}\},s)=C(\{y,y^{\prime}\},s^{\prime})$. ∎
Proof of Theorem Theorem.
@@
Suppose that $|Y|\geq 3$, that $C(\cdot,s)$ satisfies definition 2.2, and also by contradiction that it satisfies all of definitions 2.1, 2.3, and 2.4. We will prove the result in a series of steps, following the basic structure of Fey (2014). To begin, note that by Lemma 3.2 $C(\cdot,s)$ satisfies definition 3.1. Say that $j$ is decisive for $y$ over $y^{\prime}$ if $s_{j}(\cdot,y)>s_{j}(\cdot,y^{\prime})$ implies that $y\in C(\{y,y^{\prime}\},s)$.Take any $y,y^{\prime}\in Y$, $y\neq y^{\prime}$.
Step 1: identify a critical model $\mathbf{j^{*}}$. Consider the following two ensemble scores $s,s^{\prime}\in S(Y)$:
$\begin{array}[]{llll}s_{1}(\cdot,y)&>s_{1}(\cdot,y^{\prime})&>\ldots\\ \vdots&\vdots&\ddots\\ s_{m}(\cdot,y)&>s_{m}(\cdot,y^{\prime})&>\ldots\end{array}$
@@
$\begin{array}[]{llll}s^{\prime}_{1}(\cdot,y^{\prime})&>s^{\prime}_{1}(\cdot,y)&>\ldots\\ \vdots&\vdots&\ddots\\ s^{\prime}_{m}(\cdot,y^{\prime})&>s^{\prime}_{m}(\cdot,y)&>\ldots\end{array}$
By ensemble unanimity, $y=C(\{y,y^{\prime}\},s)$ and $y^{\prime}=C(\{y,y^{\prime}\},s^{\prime})$. Now, consider a set of scores $s^{j}$ such that:
$\begin{array}[]{llll}s^{j}_{1}(\cdot,y^{\prime})&>s^{j}_{1}(\cdot,y)&>\ldots\\ \vdots&\vdots&\ddots\\ s^{j}_{j}(\cdot,y^{\prime})&>s^{j}_{j}(\cdot,y)&>\ldots\\ s^{j}_{j+1}(\cdot,y)&>s^{j}_{j+1}(\cdot,y^{\prime})&>\ldots\\ \vdots&\vdots&\ddots\\ s^{j}_{m}(\cdot,y)&>s^{j}_{m}(\cdot,y^{\prime})&>\ldots\end{array}$
Define $j^{*}$ such that $y\in C(\{y,y^{\prime}\},s^{j^{*}-1})$ and $y^{\prime}\in C(\{y,y^{\prime}\},s^{j^{*}})$.
@@
Step 2: for all $\mathbf{y^{\prime\prime}\neq y,y^{\prime}}$ we have $\mathbf{j^{*}}$ decisive for $\mathbf{y^{\prime}}$ over $\mathbf{y^{\prime\prime}}$. To see this, take any $y^{\prime\prime}$ and consider $s$ such that:
$\begin{array}[]{lllll}s_{1}(\cdot,y^{\prime})&>s_{1}(\cdot,y^{\prime\prime})&>s_{1}(\cdot,y)&\ldots\\ \vdots&\vdots&\vdots&\ddots\\ s_{j^{*}-1}(\cdot,y^{\prime})&>s_{j^{*}-1}(\cdot,y^{\prime\prime})&>s_{j^{*}-1}(\cdot,y)&\ldots\\ s_{j^{*}}(\cdot,y)&>s_{j^{*}}(\cdot,y^{\prime})&>s_{j^{*}}(\cdot,y^{\prime\prime})&\ldots\\ \vdots&\vdots&\vdots&\ddots\\ s_{m}(\cdot,y)&>s_{m}(\cdot,y^{\prime})&>s_{m}(\cdot,y^{\prime\prime})&\ldots\\ \end{array}$ (9)
@@
Note that $s_{j}(\cdot,y)>s_{j}(\cdot,y^{\prime})$ if and only if $s^{j^{*}-1}_{j}(\cdot,y)>s^{j^{*}-1}_{j}(\cdot,y^{\prime})$ for all $j$ and so by Lemma 3.3 $y\in C(\{y,y^{\prime}\},s)$. By ensemble unanimity $y^{\prime}=C(\{y^{\prime},y^{\prime\prime}\},s)$. Thus by Lemma 5.1 $y=C(\{y,y^{\prime\prime}\},s)$. Next consider the set of scores $S^{*}\subseteq S(Y)$ such that for all $s^{*}\in S^{*}$:
$\begin{array}[]{lll}s^{*}_{1}(y^{\prime})\backslash s^{*}_{1}(y^{\prime\prime})&>s^{*}_{1}(y)&>\ldots\\ \vdots&\vdots&\ddots\\ s^{*}_{j^{*}}(\cdot,y^{\prime})&>s^{*}_{j^{*}}(\cdot,y)&>s^{*}_{j^{*}}(\cdot,y^{\prime\prime})\ldots\\ s^{*}_{j^{*}+1}(\cdot,y)&>s^{*}_{j^{*}+1}(\cdot,y^{\prime})&>\ldots\\ &\hskip 19.91692pt\backslash s^{*}_{j^{*}+1}(\cdot,y^{\prime\prime})&\\ \vdots&\vdots&\ddots\\ s^{*}_{m}(\cdot,y)&>s^{*}_{m}(\cdot,y^{\prime})&>\ldots\\ &\hskip 19.91692pt\backslash s^{*}_{m}(\cdot,y^{\prime\prime})&\end{array}$ (18)

### v5

@@ VI Conclusion
In this note we have analyzed a set of four desirable properties on ensemble information use and prediction stability and shown that at least one must be violated in any machine learning ensemble, establishing a fundamental tradeoff between information use and prediction instability. Any ensemble that uses the insights of multiple models will make predictions or choices in a counterintuitive or unstable fashion at least some of the time. The risk of prediction instability is created by aggregation of multiple models and heightened by the presence of high capacity individual models within the ensemble. Since three of these four properties are fairly easy to satisfy the last–model choice reversal–must fail in many common use cases thus introducing ‘discontinuities’ into the predictions and decisions made by ensembles. Small changes in underlying model scores can lead to large, substantive changes in ensemble predictions and choices; the ensemble can change its prediction or choice even if none of the underlying models do and there is no guarantee that similar feature space inputs will be treated similarly. Finally, consistency of the individual models in the ensemble implies that it is possible to find aggregation procedures that eliminate the prediction or choice instability analyzed here in asymptotic conditions.
@@
We observe that theorem III.4 as presented here shows that all ensembles will behave badly some of the time with respect to the prediction stability properties we have analyzed. It strikes us that the requirements on the model scoring function range and the surjectivity of $s(x)$ are likely much stronger assumptions in regression problems than in classification problems; though theorem III.4 holds in a technical sense in regression we speculate that the risk of running into model choice reversal violations is lower in regression problems. It also seems intuitive that the risk of prediction or choice instability depends on the number of models in the underlying ensemble and on the relationship between the size of the data sample used to train the ensemble and the population it is drawn from. Accordingly we aim to extend the analysis here to investigate asymptotic ensemble behavior when the number of models in the ensemble grows large. We conjecture that the Borda count method of aggregation may satisfy definitions II.4, II.5, and II.6 with a slight weakening of model choice reversal in the asymptotic case of many models. Finally, we conjecture that the argument developed here would apply to using a data sampling process like bootstrapping or $k$-fold cross validation to choose among multiple model architectures during training.
Acknowledgment
@@
This work was possible by the exceptional generosity of the PPC Foundation, Inc. and through support by my colleagues Christopher Taylor and Sheila Ross. Additionally, we would like to thank RJ Nowling, David Andrzejewski, and the audience of the Medical College of Wisconsin SEAWINDS Data Science Symposium for helpful comments, suggestions, and feedback. Finally, we would like to thank Ernesto Guerra Vallejos for suggestions leading to theorem V.2.
@@
Lemma VI.1.
If $C(\cdot)$ satisfies transitivity then:
@@
    $y\in C(\{y,y^{\prime}\},z)$ and $y^{\prime}=C(\{y^{\prime},y^{\prime\prime}\},z)$ implies that $y=C(\{y,y^{\prime\prime}\},z)$;
@@
    $y=C(\{y,y^{\prime}\},z)$ and $y^{\prime}\in C(\{y^{\prime},y^{\prime\prime}\},z)$ implies that $y=C(\{y,y^{\prime\prime}\},z)$;
@@
    $y=C(\{y,y^{\prime}\},z)$ and $y^{\prime}=C(\{y^{\prime},y^{\prime\prime}\},z)$ implies that $y=C(\{y,y^{\prime\prime}\},z)$.
@@
Suppose that $C(\cdot)$ satisfies transitivity and take any $z\in S(Y)$. To see part 1) assume $y\in C(\{y,y^{\prime}\},z)$ and $\{y^{\prime}\}=C(\{y^{\prime},y^{\prime\prime}\},z)$. Then by transitivity $y\in C(\{y,y^{\prime\prime}\},z)$. Suppose that $y^{\prime\prime}\in C(\{y,y^{\prime\prime}\},z)$. But then by transitivity $y^{\prime\prime}\in C(\{y^{\prime},y^{\prime\prime}\},z)$, a contradiction. To see part 2) assume $\{y\}=C(\{y,y^{\prime}\},z)$ and $y^{\prime}\in C(\{y^{\prime},y^{\prime\prime}\},z)$. Then by transitivity $y\in C(\{y,y^{\prime\prime}\},z)$. Suppose that $y^{\prime\prime}\in C(\{y,y^{\prime\prime}\},z)$. But then by transitivity $y^{\prime}\in C(\{y,y^{\prime}\},z)$, a contradiction. To see, part 3) assume that $\{y\}=C(\{y,y^{\prime}\},z)$ and $\{y^{\prime}\}=C(\{y^{\prime},y^{\prime\prime}\},z)$ but that $y^{\prime\prime}\in C(\{y,y^{\prime\prime}\},z)$. Then by part 1) we have $y^{\prime\prime}\in C(\{y^{\prime},y^{\prime\prime}\},z)$, a contradiction. ∎
Proof of Lemma III.2.
@@
Sufficiency. Suppose that $C(\cdot)$ is transitive. Take any $z\in S(Y)$, any $Y_{1}^{*},Y_{2}^{*}\subseteq Y$ such that $Y_{1}^{*}\subseteq Y_{2}^{*}$. Suppose that $Y_{1}^{*}\bigcap C(Y_{2}^{*},z)\neq\emptyset$. Take any $y\in C(Y_{1}^{*},z)$; we wish to show that $y\in Y_{1}^{*}\bigcap C(Y_{2}^{*},z)$. Clearly $y\in Y_{1}^{*}$. Suppose that $y\notin C(Y_{2}^{*},z)$. Take any $y^{\prime}\in Y_{1}^{*}\bigcap C(Y_{2}^{*},z)$. Then $y^{\prime}\in C(Y_{2}^{*},z)$ implies that $y^{\prime}\in C(,z)$, and $y\in C(Y_{1}^{*},z)$ implies that $\{y,y^{\prime}\}=C(\{y,y^{\prime}\},z)$. Since $y\notin C(Y_{2}^{*},z)$ it follows that there exists $y^{\prime\prime}\in Y_{2}^{*}$ such that $\{y^{\prime\prime}\}=C(\{y,y^{\prime\prime}\},z)$. But then by transitivity and Lemma VI.1 it must be that $\{y^{\prime\prime}\}=C(\{y^{\prime},y^{\prime\prime}\},z)$, a contradiction. Next take any $y\in Y_{1}^{*}\bigcap C(Y_{2}^{*},z)$. Then there does not exist $y^{\prime}\in Y_{2}^{*}$ such that $\{y^{\prime}\}=C(\{y,y^{\prime}\},z)$. Thus there does not exist $y^{\prime}\in Y_{1}^{*}$ such that $\{y^{\prime}\}=C(\{y,y^{\prime}\},z)$ and so $y\in C(Y_{1}^{*},z)$. It follows that $C(Y_{1}^{*},z)=Y_{1}^{*}\bigcap C(Y_{2}^{*},z)$ and so definition II.4 holds.
@@
Necessity. Suppose that $C(\cdot)$ satisfies definition II.4 but that transitivity does not hold. Then there exists $y,y^{\prime},y^{\prime\prime}$ such that $y\in C(\{y,y^{\prime}\},z)$ and $y^{\prime}\in C(\{y^{\prime},y^{\prime\prime}\},z)$, yet $\{y^{\prime\prime}\}=C(\{y,y^{\prime\prime}\},z)$. There are four cases to consider.
@@
Case 1: $\{y,y^{\prime}\}=C(\{y,y^{\prime}\},z)$, $\{y^{\prime},y^{\prime\prime}\}=C(\{y^{\prime},y^{\prime\prime}\},z)$, and $\{y^{\prime\prime}\}=C(\{y,y^{\prime\prime}\},z)$. By definition of $C(\cdot)$ we have $C(\{y,y^{\prime},y^{\prime\prime}\},z)=\{y^{\prime},y^{\prime\prime}\}$. But then we have $\{y,y^{\prime}\}\subseteq\{y,y^{\prime},y^{\prime\prime}\}$ and $\{y,y^{\prime}\}\bigcap C(\{y,y^{\prime},y^{\prime\prime}\},z)\neq\emptyset$ and yet $C(\{y,y^{\prime}\},z)=\{y,y^{\prime}\}\neq\{y^{\prime}\}=\{y,y^{\prime}\}\bigcap C(\{y,y^{\prime},y^{\prime\prime}\},z)$, a violation of definition II.4 and so a contradiction.
@@
Case 2: $\{y,y^{\prime}\}=C(\{y,y^{\prime}\},z)$, $\{y^{\prime}\}=C(\{y^{\prime},y^{\prime\prime}\},z)$, and $\{y^{\prime\prime}\}=C(\{y,y^{\prime\prime}\},z)$. By definition of $C(\cdot)$ we have $C(\{y,y^{\prime},y^{\prime\prime}\},z)=\{y^{\prime}\}$. But then we have $\{y,y^{\prime}\}\subseteq\{y,y^{\prime},y^{\prime\prime}\}$ and $\{y,y^{\prime}\}\bigcap C(\{y,y^{\prime},y^{\prime\prime}\},z)\neq\emptyset$ and yet $C(\{y,y^{\prime}\},z)=\{y,y^{\prime}\}\neq\{y^{\prime}\}=\{y,y^{\prime}\}\bigcap C(\{y,y^{\prime},y^{\prime\prime}\},z)$, a violation of definition II.4 and so a contradiction.
@@
Case 3: $\{y\}=C(\{y,y^{\prime}\},z)$, $\{y^{\prime},y^{\prime\prime}\}=C(\{y^{\prime},y^{\prime\prime}\},z)$, and $\{y^{\prime\prime}\}=C(\{y,y^{\prime\prime}\},z)$. By definition of $C(\cdot)$ we have $C(\{y,y^{\prime},y^{\prime\prime}\},z)=\{y^{\prime\prime}\}$. But then we have $\{y^{\prime},y^{\prime\prime}\}\subseteq\{y,y^{\prime},y^{\prime\prime}\}$ and $\{y^{\prime},y^{\prime\prime}\}\bigcap C(\{y,y^{\prime},y^{\prime\prime}\},z)\neq\emptyset$ and yet $C(\{y^{\prime},y^{\prime\prime}\},z)=\{y^{\prime},y^{\prime\prime}\}\neq\{y^{\prime\prime}\}=\{y^{\prime},y^{\prime\prime}\}\bigcap C(\{y,y^{\prime},y^{\prime\prime}\},z)$, a violation of definition II.4 and so a contradiction.
Case 4: $\{y\}=C(\{y,y^{\prime}\},z)$, $\{y^{\prime}\}=C(\{y^{\prime},y^{\prime\prime}\},z)$, and $\{y^{\prime\prime}\}=C(\{y,y^{\prime\prime}\},z)$. But then $C(\{y,y^{\prime},y^{\prime\prime}\},z)=\emptyset$, a contradiction. ∎
Proof of Lemma III.3.
@@
Take any $z,z^{\prime}\in S(Y)$. It suffices to consider two cases. First suppose that $C(\{y,y^{\prime}\},z)=\{y,y^{\prime}\}$ but that without loss of generality $C(\{y,y^{\prime}\},z^{\prime})=\{y\}$. Then there is an immediate contradiction of model choice reversal, since there does not exist $j$ such that $z_{j,d(y^{\prime})}>z_{j,d(y)}$ and $z^{\prime}_{j,d(y^{\prime})}<z^{\prime}_{j,d(y)}$, and so the claim holds. Next suppose without loss of generality that $C(\{y,y^{\prime}\},z)=\{y\}$ but that $y^{\prime}\in C(\{y,y^{\prime}\},z^{\prime})$. But then again there is an immediate contradiction of model choice reversal, since there does not exist $j$ such that $z_{j,d(y^{\prime})}<z_{j,d(y)}$ and $z^{\prime}_{d,d(y^{\prime})}>z^{\prime}_{j,d(y)}$ and so the claim holds. Thus $C(\{y,y^{\prime}\},z)=C(\{y,y^{\prime}\},z^{\prime})$. ∎
Proof of theorem III.4.
@@
Suppose that $|Y|\geq 3$ and that by contradiction $C(\cdot)$ satisfies definitions II.2, II.5, II.4, II.6, and II.7. We will prove (1) in a series of steps, following the basic structure of Fey (2014). To begin, note that by Lemma III.2 $C(\cdot)$ satisfies definition III.1. Say that $j$ is decisive for $y$ over $y^{\prime}$ if $z_{j,d(y)}>z_{j,d(y^{\prime})}$ implies that $\{y\}=C(\{y,y^{\prime}\},z)$. Take any $y,y^{\prime}\in Y$, $y\neq y^{\prime}$. Finally, note that all $z$ postulated below are available in $S(Y)$ by $1,\ldots,m$ being unconstrained learners.
Step 1: identify a critical model $\mathbf{j^{*}}$. Consider the following $s,s^{\prime}\in S(Y)$:
$z_{1,d(y)}$ $>z_{1,d(y^{\prime})}>\ldots$
$\vdots$
$z_{m,d(y)}$ $>z_{m,d(y^{\prime})}>\ldots$
@@
$z^{\prime}_{1,d(y^{\prime})}$ $>z^{\prime}_{1,d(y)}>\ldots$
$\vdots$
$z^{\prime}_{m,d(y^{\prime})}$ $>z^{\prime}_{m,d(y)}>\ldots$
By ensemble unanimity, $\{y\}=C(\{y,y^{\prime}\},z)$ and $\{y^{\prime}\}=C(\{y,y^{\prime}\},z^{\prime})$. Now, consider $z^{j}$ such that:
$z^{j}_{1,d(y^{\prime})}$ $>z^{j}_{1,d(y)}>\ldots$
$\vdots$
$z^{j}_{j,d(y^{\prime})}$ $>z^{j}_{j,d(y)}>\ldots$
$z^{j}_{j+1,d(y)}$ $>z^{j}_{j+1,d(y^{\prime})}>\ldots$
$\vdots$
$z^{j}_{m,d(y)}$ $>z^{j}_{m,d(y^{\prime})}>\ldots$
Define $j^{*}$ such that $y\in C(\{y,y^{\prime}\},z^{j^{*}-1})$ and $y^{\prime}\in C(\{y,y^{\prime}\},z^{j^{*}})$.
@@
Step 2: for all $\mathbf{y^{\prime\prime}\neq y,y^{\prime}}$ we have $\mathbf{j^{*}}$ decisive for $\mathbf{y^{\prime}}$ over $\mathbf{y^{\prime\prime}}$. To see this, take any $y^{\prime\prime}$ and consider $z\in S(Y)$ such that:
$z_{1,d(y^{\prime})}$ $>z_{1,d(y^{\prime\prime})}>z_{1,d(y)}\ldots$
$\vdots$
$z_{j^{*}-1,d(y^{\prime})}$ $>z_{j^{*}-1,d(y^{\prime\prime})}>z_{j^{*}-1,d(y)}\ldots$
$z_{j^{*},d(y)}$ $>z_{j^{*},d(y^{\prime})}>z_{j^{*},d(y^{\prime\prime})}\ldots$
$\vdots$
$z_{m,d(y)}$ $>z_{m,d(y^{\prime})}>z_{m,d(y^{\prime\prime})}\ldots\ldots$ (4)
@@
Note that $z_{j,d(y)}>z_{j,d(y^{\prime})}$ if and only if $z_{j^{*}-1,d(y)}>z_{j^{*}-1,d(y^{\prime})}$ for all $j$ and so by Lemma III.3 $y\in C(\{y,y^{\prime}\},z)$. By ensemble unanimity $\{y^{\prime}\}=C(\{y^{\prime},y^{\prime\prime}\},z)$. Thus by Lemma VI.1 $\{y\}=C(\{y,y^{\prime\prime}\},z)$. Next consider the set of scores $S^{*}\subseteq S(Y)$ such that for all $z^{*}\in S^{*}$:
$z^{*}_{1,d(y^{\prime})}\backslash z^{*}_{1,d(y^{\prime\prime})}$ $>z^{*}_{1,d(y)}>\ldots$
$\vdots$
$z^{*}_{j^{*},d(y^{\prime})}$ $>z^{*}_{j^{*},d(y)}>z^{*}_{j^{*},d(y^{\prime\prime})}\ldots$
$z^{*}_{j^{*}+1,d(y)}$ $>z^{*}_{j^{*}+1,d(y^{\prime})}\backslash z^{*}_{j^{*}+1,d(y^{\prime\prime})}>\ldots$
$\vdots$
$z^{*}_{m,d(y)}$ $>z^{*}_{m,d(y^{\prime})}\backslash z^{*}_{m,d(y^{\prime\prime})}>\ldots$ (5)